### v4

@@ 5.2   Weighting scheme
The previous subsection provides the method to estimate each $CLATT_{e,e+l}$ in staggered DID-IV designs. In some applications, each $CLATT_{e,e+l}$ is of intrinsic interest if researchers are interested in the treatment effect heterogeneity across cohort $e$ and relative period $l$. However, researchers may also want to estimate summary causal parameters that capture the overall effect of a treatment on an outcome. In staggered DID-IV designs, each $CLATT_{e,e+l}$ depends not only on cohort $e$, relative period $l$, but also on the compliers $CM_{e,e+l}$. In this case, each $CLATT_{e,e+l}$ may be difficult to interpret when the number of cohorts and the relative period are fairly large, and when the members of compliers $CM_{e,e+l}$ in cohort $e$ vary and overlap over time.
In this section, we explain how one can construct the summary causal parameters from each CLATT in staggered DID-IV designs, based on the weighting scheme proposed by Callaway and Sant’Anna (2021).
Specifically, we consider the following weighting scheme as in Callaway and Sant’Anna (2021):
@@
If treatment is binary, each $CETT_{e,l}^{1}$ is equal to the share of the compliers in cohort $e$ in period $e+l$:
@@
where $CATT(\tilde{e},t)$ is the cohort specific average treatment effect on the treated at time $t$ in cohort $\tilde{e}$ (see, e.g., Callaway and Sant’Anna (2021), Sun and Abraham (2021)). Callaway and Sant’Anna (2021) construct the summary measure $\theta_{sel}(\tilde{e})$ by equally weighting each $CATT(\tilde{e},t)$ across all post-treatment periods because all the units in cohort $e$ adopt the treatment after they start being exposed to it.
@@
Notes: This table represents the specific expressions for the weights on each $CLATT(e,t)$ or each $CACRT(e,t)$ in a variety of summary causal parameters. Each target parameter corresponds with that of Callaway and Sant’Anna (2021). IV is added to associate each parameter with staggered DID-IV designs.
@@
$\theta_{sel}(\tilde{e})^{IV}=\sum_{t=\tilde{e}}^{T}\frac{CAET^{1}_{\tilde{e},t}}{\sum_{t=\tilde{e}}^{T}CAET^{1}_{\tilde{e},t}}CLATT_{\tilde{e},t},$
where the weights are non-negative and sum to one by construction. This parameter summarizes each $CLATT_{\tilde{e},t}$ in cohort $\tilde{e}$ across all post-exposed periods and the weight assigned to each $CLATT_{\tilde{e},t}$ reflects the relative share of the compliers $CM_{\tilde{e},t}$ in period $t$ during the periods after the initial exposure to the instrument in cohort $\tilde{e}$. This weighting scheme would be reasonable in that it is designed to be larger in the period when the proportion of the compliers is relatively higher in cohort $\tilde{e}$.
By similar arguments, we can specify the weighting functions for various summary measures in staggered DID-IV designs, which correspond with those in Callaway and Sant’Anna (2021). In Table 7, we summarize the specific expressions for the weights assigned to each $CLATT_{\tilde{e},t}$ in each summary causal parameter. Here, we define each causal parameter by the notation analogous to that of Callaway and Sant’Anna (2021) and superscript IV to associate each parameter with staggered DID-IV designs. Note that our proposed weighting functions are the same under non-binary, ordered treatment settings, in which our target parameter is the $CACRT_{\tilde{e},t}$.
@@
where $\hat{w}(e,t)$ is the sample analog of each ${w}(e,t)$ and $\widehat{CLATT}_{e,t}$ is the consistent estimator for each $CLATT_{e,t}$, obtained from the stacked IV regression in section 5.1. Each $\hat{w}(e,t)$ is the regular asymptotically linear estimator:
@@
5.3   Pre-trend test
We now describe how one can conduct pre-trend tests to check the validity of the parallel trends assumption in the outcome and the treatment.
Suppose that data are available for period $t=-1$; one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
@@ 5.3   Pre-trend test
$\iff$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=0]=E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=1].$ (4)
These tests can be generalized to multiple pre-exposed periods settings by using the pre-trend tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al. (2021)); we can apply these tools to the first stage and the reduced form respectively to confirm the plausibility of Assumption 12 and 13.
Remark 3.
When researchers implement the staggered DID-IV design via TWFEIV regressions, they often use the dynamic specifications in the first stage and the reduced form:
$Y_{i,t}=\mu_{i.}+\delta_{t.}+\sum_{l}\alpha_{l}\mathbf{1}\{t-E_{i}=l\}+\epsilon_{i,t},$ (5)
$D_{i,t}=\gamma_{i.}+\zeta_{t.}+\sum_{l}\beta_{l}\mathbf{1}\{t-E_{i}=l\}+\eta_{i,t},$ (6)
where $E_{i}$ represents the initial exposure date. They plot the coefficients $\alpha_{l}$ in equation (5) and $\beta_{l}$ in equation (6) in each relative period $l$, and test whether the coefficients before the initial exposure to the instrument are significantly different from zero or not (e.g., Akerman et al. (2015) and Bhuller et al. (2013)).
However, such tests are problematic if the effects of the instrument on the treatment and outcome are heterogeneous. Sun and Abraham (2021) shows that in staggered DID designs, the coefficient for pre-treatment periods on event study regressions can be different from zero even when the no anticipation assumption and parallel trends assumption in the untreated outcome are plausible. Because we can view that $Y_{i,t}$ and $D_{i,t}$ are the outcomes and $Z_{i,t}$ is the binary absorbing treatment in equations (5)-(6), it is inappropriate to check whether the coefficients of pre-exposed periods are statistically different from zero or not as a way of testing Assumption 12 and Assumption 13.
@@ 6   Application
Oreopoulos (2006) estimates returns to schooling using a major education reform in the UK that increased the years of compulsory schooling from 14 to 15. Specifically, Oreopoulos (2006) exploit variation resulting from the different timing of implementation of school reforms between Britain (England, Scotland, and Wales) and Northern Ireland as an instrument for education attainment: the school-leaving age increased in Britain in 1947, but was not implemented until 1957 in Northern Ireland. The data are a sample of individuals in Britain and Northern Ireland, who were aged $14$ between $1936$ and $1965$, constructed from combining the series of U.K. General Household Surveys between $1984$ and $2006$; see Oreopoulos (2006), Oreopoulos (2008) for details.
In the first part of his analysis, Oreopoulos (2006) adopts regression discontinuity designs (RDD) and analyzes the data sets in Britain and Northern Ireland separately. Due to the imprecision of their standard errors [^8] Kolesár and Rothe (2018) consider inference in regression discontinuity designs with a discrete running variable and apply their theory to Oreopoulos (2006)., Oreopoulos (2006) then moves to ”a difference-in-differences and instrumental-variables analysis by combining the two sets of U.K. data”.
Specifically, Oreopoulos (2006) (more precisely, Oreopoulos (2008)) runs the following two-way fixed effects instrumental variable regression with the education reform as an excluded instrument for education attainment.
@@
The staggered introduction of the school reform can be viewed as a natural experiment, but is not randomized across regions in reality; Oreopoulos (2006) notes that the reform was implemented with political support, taking into account costs and benefit. This indicates that Oreopoulos (2006) implicitly relies on a staggered DID-IV identification strategy instead of the IV design. Indeed, Oreopoulos (2006) presents the corresponding plots of British and Northern Irish average education attainment and average log earnings by cohort to illustrate the evolution of these variables before and after the policy shock.
In the author’s setting, however, the effect of the school reform on education attainment and log annual earnings can differ between Britain and North Ireland, potentially varying across cohorts. That the author was conscious of the heterogeneous treatment effects is clear from the title of ”Estimating Average and Local Average Treatment Effects of Education when Compulsory Schooling Laws Really Matter”. The TWFEIV estimator, therefore, can potentially fail to aggregate the treatment effects, yielding a misleading conclusion under staggered DID-IV designs.
In this section, we first assess the identifying assumptions in staggered DID-IV designs implicitly imposed by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. Finally, we estimate the target parameter and its summary measure by employing our proposed method and weighting scheme.
@@ 6.1   Assessing the identifying assumptions in staggered DID-IV designs
We discuss the validity of the staggered DID-IV identification strategy in Oreopoulos (2006). In the author’s setting, our target parameter is the cohort specific average causal response on the treated (CACRT), as education attainment is a non-binary, ordered treatment. We can identify each CACRT and its summary measure if the underlying DID-IV identifying assumptions seem plausible.
@@
It requires that the expectation of log annual earnings would have followed the same evolution between England and North Ireland across cohorts in the absence of the school reform. Panel (b) in Figure 2 plots the results of the interacted two-way fixed effects regression in the reduced form along with a $95\%$ pointwise confidence interval. The pre-exposed estimates seems consistent with Assumption 13: though an upward pre-trend exists, all the estimates before the initial exposure to the policy change are not significantly different from zero.
Figure 2 also sheds light on the dynamic effects of the policy reform on education attainment and log annual earnings in post-exposed periods. In Panel (a), the estimated increase in years of schooling after the reform ranges from $0.56$ to $0.87$, and all estimates are statistically significant. In Panel (b), the estimated increase in log annual earnings after the reform varies from $8\%$ to $25\%$, and $5$ out of $10$ estimates are statistically significant.
Note that the post-exposed estimates in Panel (b) do not capture each CACRT in post-exposed periods: each estimate in the reduced form is not scaled by the corresponding estimate in the first stage. In the next section, we estimate each CACRT in post-exposed periods by employing our proposed estimation method.
@@ 6.2   Illustrating our estimation method
The result is shown in Table 8. The TWFEIV estimate is $-0.009$, not significantly different from 00, which indicates that, on the whole, the returns to schooling in the UK is nearly zero. However, this may be a misleading conclusion. We cannot interpret that the TWFEIV estimand captures the properly weighted average of all possible CACRTs if the effect of the school reform on education attainment or log annual earnings is not stable across cohorts (Miyaji (2024)).
The TWFEIV regression performs the ”bad comparisons” (c.f. Goodman-Bacon (2021)) in the author’s setting. When we run the TWFEIV regression, we treat the set of already exposed units in North Ireland as controls between $1957$ and $1965$, as both regions are already exposed to the policy shock during these periods. In Appendix D, we quantify the bias terms of the TWFEIV estimand by using Lemma 7 in Miyaji (2024), and show that the TWFEIV estimand is negatively biased in Oreopoulos (2006). Specifically, we show that all the bias terms are positive, whereas all the assigned weights are negative, which yields the downward bias for the TWFEIV estimand.
@@
We now employ the proposed method to estimate each CACRT. First, we create data sets by cohort $t$ ($t=1947,\dots,1956$). Each data set only contains units of cohort $t$ and cohort $t=1946$ in England and North Ireland. We define North Ireland as an unexposed group $U$ and discard the units aged at $14$ between $1957$ and $1965$ to avoid the bad comparisons. We then run TSLS regression in each data set. The standard error is calculated by using the influence function shown in equation (C.2) in Appendix C.
@@ 7   Conclusion
Instrumental variable strategies are pervasive in economics to estimate the effect of a treatment on an outcome. In this design, however, we commonly face two challenges: the internal and external validity of the IV estimate. To overcome at least external validity, many studies have leveraged variation occurring from the different timing of policy adoption across units as an instrument for treatment. The instrument based on these policy changes, however, generally fails to satisfy the randomization assumption in IV designs in practice, casting doubt on the internal validity of the IV estimate. Rather, the underlying identification strategy is parallel to DID designs, under which we identify the effect of the policy shock on the outcome, leveraging the variation of the exposure to that policy shock across units and over time.
@@
In this paper, we formalize the underlying identification strategy as an instrumented difference-in-differences (DID-IV) by combining the IV techniques with DID designs. We start with a simple two periods and two groups setting. In this setting, our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups. We show that the Wald-DID estimand captures the LATET under $2\times 2$ DID-IV designs. We also clarify the interpretation of the parallel trends assumption in the outcome. Specifically, we show that in DID-IV settings, time also plays the role of instrument, and the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups. Finally, we compare DID-IV to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and point out the issues inherent in Fuzzy DID designs.
@@
Next, we consider the DID-IV design in more than two time periods with units being exposed to the instrument at different times. We call this a staggered DID-IV design, and formalize the target parameter and identifying assumptions. Our target parameter in this design is the cohort specific local average treatment effects on the treated (CLATT). The identifying assumptions are the natural generalization of those in $2\times 2$ DID-IV designs.
@@
We also provide the estimation method in staggered DID-IV designs, which does not require strong restrictions on treatment effect heterogeneity. In practice, most DID-IV applications implement the staggered DID-IV design by running TWFEIV regressions. In companion paper (Miyaji (2024)), however, we show that the conventional approach is inadequate for summarizing the treatment effects if the effect of the instrument on the treatment or the outcome is not stable over time. Our estimation method carefully chooses the comparison groups and does not suffer from the bias arising from the time-varying exposed effects. We also propose the weighting scheme in staggered DID-IV designs, and explain how one can conduct pre-trend tests to assess the validity of the parallel trends assumption in the treatment and the outcome.
Finally, we illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the UK, exploiting the timing variation of the introduction of school reforms between British and North Ireland. First, we assess the underlying DID-IV identification strategy in the author’s setting and check whether the identifying assumptions are plausible. Next, we estimate a TWFEIV regression and our stacked two stage least squares regressions in each dataset. The TWFEV estimate turns out to be strictly negative, whereas our STS estimates are all positive. Finally, we estimate the summary causal parameter by exploiting our STS estimates. The estimate for the summary measure is also positive and indicates that returns to schooling are substantial in the UK.
@@
Overall, this paper provides a new econometric framework for estimating the causal effects when the treatment adoption is potentially endogenous over time, but researchers can use variation in the timing of policy adoption across units as an instrument for the treatment. To avoid the issue of using the TWFEIV estimator under this design, we also provide a reliable estimation method that is free from strong restrictions on treatment effect heterogeneity. Further developing alternative estimation methods and diagnostic tools will be a promising area for future research, facilitating the credibility of DID-IV research design in practice.
Appendix A Proofs in Section 2
@@
First we consider the denominator of the Wald-DID estimand. For the denominator, we have
$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$ (9)
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
$=$ $Pr(D_{1}((0,1))>D_{1}((0,0))|E=1)$
$=$ $Pr(CM^{Z}|E=1).$ (10)
where the second equality follows from the simple manipulation. The third equality follows from Assumption 4 (No anticipation in the first stage) and the forth equality follows from Assumption 3 (Monotonicity assumption) and Assumption 5 (Parallel trends assumption in the treatment).
Next, to show that the Wald-DID estimand equals the LATET, it suffices to show that
$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$ $=E[Y_{1}(1)-Y_{1}(0)|E=1,CM^{Z}]\cdot Pr(CM^{Z}|E=1)$
$=LATET\cdot Pr(CM^{Z}|E=1).$
For the numerator, we have
@@
The first equality follows from Assumption 1 (No carryover assumption) and Assumption 2 (Exclusion restriction). The second equality follows from the simple manipulation. The third equality follows from Assumption 4. The fourth equality follows from Assumption 6 (Parallel Trends Assumption in the outcome) and simple calculation. The fifth equality follows from the Law of Iterated Expectations and Assumption 3.
Combining the result (11) with (10), we obtain
$\frac{E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]}{E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=1]}$ $=\frac{LATET\cdot Pr(CM^{Z}|E=1)}{Pr(CM^{Z}|E=1)}$
$=LATET.$
This completes the proof. ∎
Proof of Theorem 2.
Recall that we define $\Delta_{e}$ as follows:
$\Delta_{e}\equiv E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=e].$
Then, we obtain:
$\Delta_{e}$ $=E[Y_{1}(1)-Y_{0}(1)|AT^{T},E=e]Pr(AT^{T}|E=e)$
$+E[Y_{1}(0)-Y_{0}(0)|NT^{T},E=e]Pr(NT^{T}|E=e)$
$+E[Y_{1}(1)-Y_{0}(0)|CM^{T},E=e]Pr(CM^{T}|E=e)$
$+E[Y_{1}(0)-Y_{0}(1)|DF^{T},E=e]Pr(DF^{T}|E=e)$
$=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e}.$
The first equality follows from the Law of Iterated Expectation and the second equality follows from the definition of $w_{g^{t},e}$ and $\Delta_{g^{t},e}$. Clearly, the weight $w_{g^{t},e}$ sums to one by construction.
This completes the proof.
Appendix B Extensions
B.1   Non-binary, ordered treatment
@@
Let $\lambda_{Z,j}^{t}=\mathbf{1}(D_{t}((0,z))\geq j)$ be the indicator function for $t\in\{0,1\},z\in\{0,1\}$ and $j\in\{0,\dots,J+1\}$. We note that $\lambda_{Z,0}^{t}=1$ and $\lambda_{Z,J+1}^{t}=0$ hold for all $t$ and $z$ by construction. Here, we can rewrite the observed outcome $Y_{t}$ as follows:

### v5

@@ 5.2   Weighting scheme
In this subsection, we explain how one can construct the summary causal parameters from each CLATT in staggered DID-IV designs, based on the weighting scheme proposed by Callaway and Sant’Anna (2021).
We consider the following weighting scheme as in Callaway and Sant’Anna (2021):
@@
If treatment is binary, each $CAET_{e,l}^{1}$ is equal to the share of the compliers in cohort $e$ in period $e+l$:
@@
where $CATT(\tilde{e},t)$ is the cohort specific average treatment effect on the treated at time $t$ in cohort $\tilde{e}$ (see, e.g., Callaway and Sant’Anna (2021), Sun and Abraham (2021)).
@@
Notes: This table represents the specific expressions for the weights on each $CLATT(e,t)$ or each $CACRT(e,t)$ in a variety of summary causal parameters. Each target parameter without the superscript IV is defined in Callaway and Sant’Anna (2021). The superscript IV is added to associate each parameter with staggered DID-IV designs.
@@
$\theta_{sel}(\tilde{e})^{IV}=\sum_{t=\tilde{e}}^{T}\frac{CAET^{1}_{\tilde{e},t}}{\sum_{t=\tilde{e}}^{T}CAET^{1}_{\tilde{e},t}}CLATT_{\tilde{e},t}.$
This parameter summarizes each $CLATT_{\tilde{e},t}$ in cohort $\tilde{e}$ across all post-exposed periods and the weight assigned to each $CLATT_{\tilde{e},t}$ reflects the relative share of the compliers $CM_{\tilde{e},t}$ in period $t$ during the periods after the initial exposure to the instrument in cohort $\tilde{e}$. This weighting scheme would be reasonable in that it is designed to be larger in the period when the proportion of the compliers is relatively higher in cohort $\tilde{e}$.
By similar arguments, we can specify the weighting functions for various summary measures, which correspond with those in Callaway and Sant’Anna (2021). Table 5 summarizes the specific expressions for the weights assigned to each $CLATT_{\tilde{e},t}$ in each summary causal parameter. Note that our proposed weighting functions are the same under non-binary, ordered treatment settings, in which our target parameter is the $CACRT_{\tilde{e},t}$.
@@
where $\hat{w}(e,t)$ is the sample analog of each ${w}(e,t)$ and $\widehat{CLATT}_{e,t}$ is the consistent estimator for each $CLATT_{e,t}$, obtained from the STS regression in Subsection 5.1. Each $\hat{w}(e,t)$ is the regular asymptotically linear estimator:
@@
5.3   Pre-trends test
In most DID applications, researchers assess the plausibility of the parallel trends assumption by testing for pre-treatment differences in the outcome between treatment and control groups (pre-trends test). In this section, we describe the procedure of the pre-trends test in DID-IV designs to check the validity of the parallel trends assumption in the treatment and the outcome.
Suppose that data are available for period $t=-1$. Then, one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
@@ 5.3   Pre-trends test
$\iff$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=0]$
$=$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=1].$ (2)
These tests can be generalized to multiple pre-exposed periods settings by using the pre-trends tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al. (2024)); that is, one can apply these tools to the first stage and the reduced form ($D_{i,t}$ or $Y_{i,t}$ is outcome, $Z_{i,t}$ is treatment) respectively to confirm the plausibility of Assumption 13 and 14 in pre-exposed periods.
@@ 6   Application
We illustrate the empirical relevance of our findings in the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom. We first assess the identifying assumptions in staggered DID-IV designs implicitly imposed by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. Finally, we estimate the target parameter and its summary measure by employing our proposed method and weighting scheme.
6.1   Setting
Oreopoulos (2006) estimates returns to schooling using a major education reform in the UK that increased the years of compulsory schooling from 14 to 15. Specifically, Oreopoulos (2006) exploits variation resulting from the different timing of implementation of school reforms between Britain (England, Scotland, and Wales) and Northern Ireland as an instrument for education attainment: the school-leaving age increased in Britain in 1947, but was not implemented until 1957 in Northern Ireland. [^19] In the first part of his analysis, Oreopoulos (2006) adopts regression discontinuity designs (RDD) and analyzes the data sets in Britain and Northern Ireland separately. Due to the imprecision of their standard errors, Oreopoulos (2006) then moves to “a difference-in-differences and instrumental-variables analysis by combining the two sets of U.K. data”. The data are a sample of individuals in Britain and Northern Ireland, who were aged $14$ between $1936$ and $1965$, constructed from combining the series of U.K. General Household Surveys between $1984$ and $2006$; see Oreopoulos (2006), Oreopoulos (2008) for details.
Oreopoulos (2006) (more precisely, Oreopoulos (2008)) runs the following two-way fixed effects instrumental variable regression with the education reform as an excluded instrument for education attainment.
@@ 6.1   Setting
The staggered introduction of the school reform can be viewed as a natural experiment, but is not randomized across regions in reality; Oreopoulos (2006) notes that the reform was implemented with political support, taking into account costs and benefit. This indicates that Oreopoulos (2006) implicitly relies on a staggered DID-IV identification strategy instead of exploiting the random variation of the policy change. Indeed, Oreopoulos (2006) presents the corresponding plots of British and Northern Irish average education attainment and average log earnings by cohort to illustrate the evolution of these variables before and after the policy shock.
@@ 6.2   Assessing the identifying assumptions in staggered DID-IV designs
We first discuss the validity of the staggered DID-IV identification strategy in Oreopoulos (2006). In the author’s setting, our target parameter is the cohort specific average causal response on the treated (CACRT), as education attainment is a non-binary, ordered treatment.
@@
It requires that the expectation of log annual earnings would have followed the same evolution between England and North Ireland across cohorts in the absence of the school reform. Panel (b) in Figure 1 plots the results of the interacted two-way fixed effects regression in the reduced form along with a $95\%$ pointwise confidence interval. The pre-exposed estimates seems consistent with Assumption 14: though an upward pre-trends exists, all the estimates before the initial exposure to the policy change are not significantly different from zero.
Figure 1 also sheds light on the dynamic effects of the school reform on education attainment and log annual earnings in post-exposed periods. In Panel (a), the estimated increase in years of schooling after the reform ranges from $0.56$ to $0.87$, and all estimates are statistically significant. In Panel (b), the estimated increase in log annual earnings after the reform varies from $8\%$ to $25\%$, and $5$ out of $10$ estimates are statistically significant. [^20] Note that the post-exposed estimates in Panel (b) do not capture each CACRT in post-exposed periods: each estimate in the reduced form is not scaled by the corresponding estimate in the first stage.
@@ 6.3   Illustrating our estimation method
The result is shown in Table 6. The TWFEIV estimate is $-0.009$ and not significantly different from 00. This indicates that, on the whole, the returns to schooling in the UK is nearly zero. However, this may be a misleading conclusion. We cannot interpret that the TWFEIV estimand captures the properly weighted average of all possible CACRTs if the effect of the school reform on education attainment or log annual earnings is not stable across cohorts (Miyaji (2024)).
In online Appendix Section C, we quantify the bias terms of the TWFEIV estimand by using Lemma 7 in Miyaji (2024), and show that the TWFEIV estimand is negatively biased in Oreopoulos (2006). Specifically, we show that all the bias terms are positive, while all the assigned weights are negative, which yields the downward bias for the TWFEIV estimand. [^21] When we run the TWFEIV regression in the author’s setting, we treat the set of already exposed units in North Ireland as controls between $1957$ and $1965$, as both regions are already exposed to the policy shock during these periods. This procedure performed by the TWFEIV regression is called the “bad comparisons” in the recent DID literature (c.f. Goodman-Bacon (2021)), and yields the bias in the TWFEIV estimand from the properly weighted average of all possible CACRTs.
@@
We now employ the proposed method to estimate each CACRT. First, we create data sets by cohort $t$ ($t=1947,\dots,1956$). Each data set only contains units of cohort $t$ and cohort $t=1946$ in England and North Ireland. We define North Ireland as an unexposed group $U$ and discard the units aged at $14$ between $1957$ and $1965$ to avoid the bad comparisons. We then run TSLS regression in each data set. The standard error is calculated by using the influence function shown in equation (B) in online Appendix Section LABEL:ApeC.1.
@@ 7   Conclusion
In this paper, we formalize an instrumented difference-in-differences (DID-IV). First, we consider a simple setting with two periods and two groups. In this setting, our DID-IV design mainly comprises a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups. We show that in $2\times 2$ DID-IV designs, the Wald-DID estimand captures the local average treatment effect on the treated (LATET). After establishing $2\times 2$ DID-IV designs, we compare DID-IV to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and point out the issues inherent in Fuzzy DID designs.
@@
Next, we consider the DID-IV design in more than two periods with units being exposed to the instrument at different times. We call this a staggered DID-IV design, and formalize the target parameter and identifying assumptions. Specifically, in this design, our target parameter is the cohort specific local average treatment effects on the treated (CLATT). The identifying assumptions are the natural generalization of those in $2\times 2$ DID-IV designs.
@@
We also provide the estimation method in staggered DID-IV designs that does not require strong restrictions on treatment effect heterogeneity. Our estimation method carefully chooses the comparison groups and does not suffer from the bias arising from the time-varying exposed effects. We also propose the weighting scheme in staggered DID-IV designs, and explain how one can conduct pre-trends tests to assess the validity of the parallel trends assumptions in DID-IV designs.
Finally, we illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting the timing variation of the introduction of school reforms between British and North Ireland. In this application, the TWFEV regression, the conventional approach to implement a staggered DID-IV design, yields a negative estimate. By contrast, our STS regression and weighting scheme indicate the substantial gain from schooling. This empirical application illustrates that the estimation method matters in staggered DID-IV designs in practice.
@@
Overall, this paper provides a new econometric framework for estimating the causal effects when the treatment adoption is potentially endogenous over time, but researchers can exploit variation in policy adoption timing as an instrument for treatment. To avoid the issue of using the TWFEIV estimator, we also provide a reliable estimation method that is free from strong restrictions on treatment effect heterogeneity. Further developing alternative estimation methods and diagnostic tools will be a promising area for future research, facilitating the credibility of DID-IV design in practice.
Appendix
@@
First we consider the numerator of the Wald-DID estimand. For the numerator, we have
@@
The first equality follows from Assumption 1 and Assumption 2. The second equality follows from the simple manipulation. The third equality follows from Assumption 4. The fourth equality follows from Assumption 6 and simple calculation. The fifth equality follows from the Law of Iterated Expectations and Assumption 3.
For the denominator, we have
$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
$=$ $Pr(CM^{Z}|E=1).$ (6)
where the second equality follows from the simple manipulation. The third equality follows from Assumption 4 and the forth equality follows from Assumption 3 and Assumption 5.
Combining the result (5) with (6), we obtain the desirable result. ∎
@@
First, we prove the consistency of our SLS estimator. We assume that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are well defined. More precisely, we assume that cohort $e$ is non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$. This assumption guarantees that each Wald-DID estimator $\widehat{CLATT}_{e,e+l}$ is also well defined.
First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. By the Law of Large Numbers and continuous mapping theorem, we have
[…]
Next, we show that the right hand side of (7) is equal to $CLATT_{e,e+l}$. For the numerator, we have
$E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]$ (8)
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]$
The third equality follows from Assumption 9, 10 and 12. The fourth equality follows from Assumption 14 and the fifth follows from Assumption 11. For the denominator, by similar calculations,
$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]$
$=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]+E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]$ (10)
$-\sum_{e\in U}E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
$=Pr(CM_{e,e+l}|E_{i}=e),$ (11)
where the final equality follows from Assumption 11 and Assumption 13. Combining (9) with (11), we obtain the desirable result.
Next, we prove that our STS estimator is asymptotically normal, deriving its influence function. We define $\theta_{e,l}$ to be
$\equiv\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}.$
The following fact, also found in de Chaisemartin and D’Haultfœuille (2018), is repeatedly used in the derivation of the influence function.
Fact 1.
If
we have
Then, by some calculations, we obtain
$\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}\psi_{i,e,l}+o_{p}(1).$
The influence function $\psi_{i,e,l}$ is:
$\psi_{i,e,l}$
Therefore, the asymptotic variance $V(\psi_{i,e,l})$ is
$V(\psi_{i,e,l})$ $=\frac{1}{(\pi_{e,l}^{1}-\pi_{e,l}^{2})^{2}}\Bigg{(}\frac{E\big{[}\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=e\}]^{2}}$
$+\frac{E\big{[}\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=U\}]^{2}}\Bigg{)}.$
Appendix A Extensions
This section considers the non-binary, ordered treatment and repeated cross section settings in DID-IV designs. It also includes the case where we introduce the treatment path in $2\times 2$ DID-IV designs.
A.1   Non-binary, ordered treatment
In this subsection, we present the identification results in DID-IV designs with non-binary, ordered treatment. Theorem 4 below shows that if we have a non-binary, ordered treatment, the Wald-DID estimand is equal to the ACRT under Assumptions 1-7, which replace the binary treatment with a non-binary one.
Theorem 4.
If Assumptions 1-7 hold, the Wald-DID estimand $W_{DID}$ is equal to the ACRT.
Proof.
Let $\lambda_{Z,j}^{t}=\mathbf{1}(D_{t}((0,z))\geq j)$ be the indicator function for $t\in\{0,1\},z\in\{0,1\}$ and $j\in\{0,\dots,J+1\}$. Note that $\lambda_{Z,0}^{t}=1$ and $\lambda_{Z,J+1}^{t}=0$ hold for all $t$ and $z$ by construction. Here, we can rewrite the observed outcome $Y_{t}$ as follows: